OpenStudy (eric_d):

Given that f(x)=x^3+kx^2-2x+1 has reamainder k when it is divided by (x-k),find possible values of k @Callisto

OpenStudy (callisto):

Use the remainder theorem. When f(x) is divided by x=m and have a remainder n, we have f(m) = n In this case, f(x) is divided by k and the remainder is k, so, what is m and n?

OpenStudy (eric_d):

f(k)=k

OpenStudy (eric_d):

k^3+k(k)^2-2k+1=k k^3+k^3-3k+1=0 2k^3-3k+1=0 like this @Callisto

OpenStudy (callisto):

Yes. Now, try factor theorem to factorize the expression on the left.

OpenStudy (eric_d):

nt sure..

OpenStudy (callisto):

By educated guess, try x=1...

OpenStudy (eric_d):

f(1)=k

OpenStudy (callisto):

Sorry... Consider f(k) = 2k^3-3k+1, what is f(1)?

OpenStudy (eric_d):

, try x=1..

OpenStudy (callisto):

Ehmmmm.... If g(k) = 2k^3-3k+1, what is g(1)?

OpenStudy (eric_d):

g(1)=0 so, (x-1) is a factor of g(k)

OpenStudy (callisto):

(k-1) is a factor of g(k), there is no x in the function g. Now, divide g(k) by k-1 to get the other factor. Can you do it?

OpenStudy (eric_d):

 not sure..